OpenStudy (anonymous):

Having a brain freeze. I am solving my problem and I have to put my reasoning but just cannot, for the life of me, remember the word that explains this step. I know its really simple and I'll feel like a noob when this gets answered lol. Help me out D: 2+cos(x+y)+cos(x+y)= 2+2cos(x+y)

OpenStudy (anonymous):

what are you solving for x and y?

hero (hero):

2+cos(x+y)+cos(x+y)= 2+2cos(x+y) cos(x+y)+cos(x+y)= 2+2cos(x+y) - 2 2cos(x+y) = 2 - 2 + 2cos(x+y) 2cos(x+y) = 2cos(x+y)

OpenStudy (anonymous):

a + a = 2a let a = cos (x + y) then cos (x + y) + cos (x + y) = 2cos (x + y)

hero (hero):

@bluepig148 , at least your solution is simpler